Ryan has the numbers 1–30 listed on individual index cards. If Ryan randomly selects 1 card, what is the probability that it will be a multiple of 3 or a multiple of 11?
Answer by ewatrrr(24785)

 
Hi,
numbers 1–30 listed on individual index cards
Includes 10 Multiples of 3: 3,6,9,12,15,18,21,24,27,30
Includes 2 Mulitples of 11: 11,22
1 card selected
P(multiple of 3 or a multiple of 11) = 10/30 + 2/30 = 12/30 = 2/5